If the rectangle is one whole, show the following:

One sixth

Three eighths

Seven quarters


1
Expert's answer
2021-06-03T12:54:14-0400

Solution.

If the rectangle is one whole, "\frac{1}{6}" is:



"\frac {3}{8}" is:



"\frac{7}{4}=1\frac{3}{4}" . The fraction "\frac{7}{4}" is more than one whole rectangle. "\frac{7}{4}" is: